### v7

# Deep learning architecture for decrypting information on the event horizon

Abstract
According to βt Hooft, to recover the invariance under the PoincarΓ© group in a holographic setting, the evolution law for the direction orthogonal to the given surface and the time evolution law must commute. The condition of commutativity assumes that the time-evolution law on the given surface and the target surface is the same. Meanwhile, the AdS/CFT correspondence implies that there exists a mapping from the black hole horizon to the conformal boundary. In this paper, it is conjectured that in a large AdS black hole in thermal equilibrium, this mapping and the time-evolution law will commute using the properties derived from the infinite volume limit. Also, we provide a computational model in the context of AdS/CFT correspondance assuming this commutativity in a holographic setting. Specifically, we will use deep learning algorithms to find a model that tells us how pieces of information on the event horizon of an AdS black hole are decrypted into conformal boundary. By combining commutativity with AdS/CFT, we might move beyond the view that the conformal boundary implements the holographic screen and find clues about how information pieces on the event horizon are transformed into information pieces on the conformal boundary.
1 Introduction
The holographic principle is considered a fundamental characteristic of quantum gravity theory. This principle started from the fact that the entropy of a black hole is proportional to the area of the horizon and was extended to the general case[1]. Considering unitarity, entropy and counting arguments, it was pointed out that the 2-dimensional surface should determine all information in the 3+1 dimension. Note that considering the limit where the two-dimensional space becomes flat, this means that there is a transform in the direction perpendicular to that flat space. Furthermore, since the ultimate theory will satisfy PoincarΓ© invariance, this transformation and the time evolution will commute with each other. The proposed cellular automaton model, in which information in a two-dimensional space determines all information in the 3+1 dimension, had only a linear structure and could not include interactions. Therefore, we aim to build a non-linear model from an AdS/CFT perspective.
The flat space-time used to construct this model was implemented as a holographic screen, and its connection with string theory was revealed[2]. For the event horizon, it can actually be mapped by following light rays. Here, the focusing theorem guarantees that there exists at least one map satisfying the holographic principle. This mapping unfolds the entropy on the black holeβs horizon onto flat surface, so black hole physics still lives on this screen. However, since it does not evolve according to Lorentzian time on a holographic screen, it is difficult to define a cellular automaton model on this screen[3]. Moreover, since the degree of freedom is not generally preserved, it is difficult to expect it to follow conventional physics. Instead, we will assume that the local nature of physics allows us to select small pieces of information from sufficiently large horizon and conformal boundary respectively, and based on this assumption, an approximation method that flattens the selected information pieces will be adopted.
In the AdS/CFT correspondence, the conformal boundary is realized as a holographic screen[4]. This correspondence guarantees the existence of a transformation that connects the event horizon and the conformal boundary. The assumption of commutativity is a very strong one, thus finding a model that satisfies this property in a typical AdS/CFT setting is inevitably very limited. Therefore, we will only consider the thermal equilibrium state of the large AdS black hole. In the context of the black hole paradox, the piece of information inscribed on the horizon can be considered to be the encrypted information of the conformal boundary[5]. From this perspective, the function that goes from the horizon to the conformal boundary will be called the decrypting function. Here, saying that this function is reversible is equivalent to saying that no information is lost in a black hole. In this paper, we propose a computational model that can find this decrypting function with minimal physical constraints.
Meanwhile, in the field of deep learning, there are three notable research results related to the general relativity, quantum field theory and AdS/CFT. Firstly, there has been a study on convolutional neural networks(CNN) in non-Euclidean space to effectively process data in non-Euclidean space such as social networks, medical information, brain imaging and computer graphics. For example, training a conventional CNN after projecting the data from the sphere onto a plane fails, but by constructing CNN that is invariant under rotation, it can be trained successfully with this data[6][7]. Furthermore, this concept can be extended to include gauge symmetry and implemented through Icosahedral CNN[8]. As the concept of symmetry was introduced into CNN, elements necessary for convolution were mapped to geometric objects, and consequently, convolution could be expressed in covariant form[9][10].
Also, there is a case of studying the AdS/CFT correspondance by mapping it to deep learning[11][12]. In this study, the bulk metric function on the black hole horizon can be trained given the data at the boundary of the quantum field theory. The philosophy is that the bulk spacetime can correspond to a deep neural network.
Finally, since gauge/gravity theory and deep learning can be expressed as a matrix model, the correspondence between them have been considered[13]. In this work, for restricted Boltzmann machine and cubic learning system, matter fields are mapped to a layer of given deep learning architecture and gauge/gravity fields are mapped to weight matrix. Furthermore, the learning process, which is the time-evolution of weight matrix, can be thought of as a process in which the effective law changes, such as spontaneous symmetry breaking.
These results imply that deep learning can be an effective computational tool even in quantum gravity. Therefore, the proposed computational model adopts a deep learnig algorithm.
2 Commutativity and AdS/CFT correspondonce
The object to be dealt with here is the AdS black hole, so the many-body problem will be considered. Additionally, for the purpose of building the computational model of AdS/CFT, we will assume that a time evolution law of microstates forming an ensemble exists as in the cellular automaton(CA) interpretation, but we will exclude fundamental discussions of quantum mechanics such as the Copenhagen interpretation[14]. Furthermore, we will assume that locality is satisfied.
The holographic principle implies the existence of an evolution law in the holographic direction. This law will be called the operator. Assuming that the invariance under the PoincarΓ© group is recovered in the ultimate holographic theory, the time evolution law and must commute[1]. Furthermore, to assume this invariance, the spaces considered must be flat. Let us first consider an approximation for flatness in typical AdS/CFT setting with block hole. In this setting, the conformal boundary is considered as the spacetime where the information on horizon is decrypted. To meet requirement for flatness, we will use the fact that the conformal boundary becomes locally flat when the black holeβs mass becomes infinitely large[15]. In other words, it has the same effect as taking an inifinite volume limit. Then the boundary becomes that is flat. Note that the physics above this conformal boundary is described by gauge theory, so it is different from the holographic screen. Also, for the event horizon, as the mass of the black hole increases, the radius of the horizon can be infinitely increased. Therefore, we can make the conformal boundary and horizon as large as we want. If we select sufficiently small information pieces from a very large surface, we can say that these pieces are approximately flat.
Additionally, for large AdS black hole, the radius of the horizon is much larger than the radius of curvature of the AdS in this limit. Therefore, if the size of the black hole becomes very large but is finite, the entropy above the conformal boundary and horizon is proportional. Moreover, we can expect the entropies to be equal at an appropriate βt Hooft parameter . Therefore, to compare the same amount of entropy, finite scale will be considered in the proposed model.
Combining the two preceding conditions, the proposed model considers a sufficiently large but finite volume of the boundary. Indeed, due to locality, it can be expected that such information pieces exist on each surface and that they will be simply connected. These selected pieces of information are related by transform. With this setting, a more general form of commutativity can be written as Equation 1, where the our assumption says that and .
Here, and are the time evolution law at the conformal boundary of AdS space and event horizon of large AdS black hole, respectively. They can be simply denoted as if .
To find a specific model that satisfies and in 1, consider the large Schwarzschild black hole solution in space. In this case, it is known that gauge theory above the conformal boundary has high temperature behavior[15]. Also, the corresponding AdS black hole horizon can reach thermal equilibrium[16]. Therefore, any piece of information on the conformal boundary and the corresponding piece of information on the event horizon can be considered as one possible microstate. Here, the change over time in the piece of information inscribed on the horizon is due to quantum fluctuations near the horizon. The Hawking temperature formula tells us the rate at which quantum fluctuations occur, and since this is expressed in coordinate time, pieces of information also change according to coordinate time. Additionally, after the black hole in AdS reaches thermal equilibrium, the emission and absorption rate must be the same, so the absorption rate also changes depending on coordinate time. Specifically, the rate at which Hawking-radiated particles are reflected at the boundary and re-absorbed has the same magnitude and opposite sign as the emissivity due to Hawking radiation[17]. Therefore, coordinate time can be given as a time parameter for the microscopic state on the event horizon same as on the conformal boundary. Also, the AdS/CFT correspondence predicts that a state corresponding to the quantum fluctuations on the horizon also exists on the conformal boundary. Furthermore, since it is expected that PoincarΓ© invariance will be restored at the Planck scale, it is conjectured that the operator and the operator for information pieces on these boundaries will commute. Note that the duality can be confirmed more clearly through deformation by adding branes to given space[18]. This might be considered a deformation where is maintained but becomes a value other than 1 in Equation 1.
Now, considering all these points in this section, as shown in Figure1 a), commutativity means that the results of time evolution at the event horizon and then decrypting to a conformal boundary are the same as those of time evolution after decrypting to a conformal boundary.
3 Deep learning for holographic system
Assuming that we know the time evolution as in CA, let us implement a computational model to find the operator using a deep learning algorithm. In other words, assuming that is given, will be trained as in Figure1 b). The spatial dimension of the event horizon corresponds to the spatial dimension of the conformal boundary. For visualization purposes, we consider here an AdS/CFT setting with one spatial dimension reduced, i.e., where the spatial dimensions of the boundary of AdS space and the horizon of AdS black hole are both 2. It is straightforward to recover that dimension in proposed models. Additionally, for simplicity, classical CA will be considered, but the full description must be quantum mechanical, and therefore the computational model should also adopt a quantum mechanical deep learning algorithm.
To implement the proposed computational model, it is crucial to note that in deep learning algorithms, updating weights occur with each batch. Consequently, labels will be generated using the intermediate state of the model updated up to the given batch point. Therefore, as illustrated in the proposed algorithm below1, the loops should be decomposed into those for epochs and those for batches.
In the above proposed algorithm, various functions are employed to facilitate the process. IntermediateModel is a model trained up to the current batch point. By placing intermediate states in a loop, learning is possible recursively. GenerateRandomBinaryTensor function generates raw random data to be used as input for the model. PreprocessData formats raw data to make it suitable for model input, ensuring it is appropriately normalized and transformed. GetBatch function extracts a specific batch from the entire dataset, defined by the batch size, and uses the given batch number and size to partition the data, returning the specified batch. LogitToPrediction converts the modelβs output, known as logits, into final predictions. This conversion is typically achieved through softmax or sigmoid functions, which transform logits into probability values that serve as the modelβs predictions. ConvertToOneHot transforms the given label data into one-hot encoded vectors, which are essential for multi-class classification tasks. Finally, TrainModel uses the provided input and target data to train the model, adjusting the model parameters to minimize the loss function throughout the training process.
For the operator assumed above, at least the following two properties must be satisfied. The first condition is the locality assumed earlier. This is satisfied if the current cell is determined by the neighboring cells in the specified bound like Conwayβs game of life. Second, roughly speaking, to preserve quantum mechanical probability it should be time-reversible. That is, the cells in the past should be able to be determined from current cells. Since two consecutive CA layers are needed to define the rule going to the future or the past, the evolution law for one time step can be split into a two-step procedure. Here, the block CA [19] will be adopted as an example of time-evolution law. The well-known block CA rules are as follows.
For edge cells after the first rule is applied, it is assumed that cells on opposite sides are grouped together to form a torus. It is known that it satisfies locality by definition and is time-reversible. Note that when we know the rules for even and odd sites like block CA, it is known that Hamiltonian can be expressed as an infinite sequence of commutators of these rules using BakerβCampbellβHausdorff formula[14].
On the other hand, we know that the block CA rule must satisfy linearity to become the time-evolution law of the quantum system. The fact that the linearity is satisfied can be seen easily by showing that the block CA rule is expressed as a matrix multiplication. This matrix is block diagonal and the elements of this block diagonal matrix are functions that, given initial conditions, become 0 or 1 over time. If the block CA were a real quantum system, each element would be a time-dependent function determined by SchrΓΆdinger equation, Hamiltonian, and initial condition.
As a dataset, random input will be used in this paper. A random dataset is defined on a event horizon, where the values 0βs and 1βs on each lattice might indicate the absence or presence of a photon at that position, respectively. After mapping information to the lattice like this example, the operator converts it to information in the corresponding conformal boundary. Although discussions on how to encode physical states into information using bits as the unit are crucial in the proposed computational model, it is not the objective of this paper.
Now, the architecture of CNN to learn operater is presented in Figure2.
Here, a convolution-deconvolution[20] architecture was used recursively. The proposed network follows a conventional deep learning architecture used in image segmentation, characterized by a convolution-max pooling-deconvolution block. The depth of these blocks is considered irrelevant to the current problem, as the correlation between the depth and the loss is not discernible critically. However, dropout layers and regularization are critically important for the problem we aim to solve. This is because, among the solutions that satisfy commutativity, there are cases where is a constant function or trivial solutions. Without these techniques, the model tends to learn trivial solutions.
However, more importantly, it is crucial to carefully select the loss function. As mentioned earlier, the proposed model tends to learn trivial solutions, so to avoid this, the loss function must be customized to penalize trivial solutions. In this model, we added a conditional statement to assign a very large loss value when a trivial solution is encountered, thereby encouraging the learning of non-trivial solutions. Nonetheless, non-trivial solutions are not necessarily physically meaningful. To obtain physically meaningful solutions, the model must incorporate the requirements derived from the holographic principle in the context of AdS/CFT. For example, among the numerous solutions that can take, the selected solution often depends on the weight initialization. For more realistic learning, it is essential to initialize the weight based on a physically correct interpretation of the weight[11].
It can be confirmed in Figure3 that our architecture is able to learn the operator.
Here, white represents a living cell and black represents a dead cell. During the inference process, the well-known glider pattern from the game of life was used as the input. However, it can be confirmed that the output after applying is the same as the output after applying , even with different inputs. Therefore, the proposed algorithm can learn a function that decrypts information pieces on a toy model of the event horizon to information pieces on a toy model of the conformal boundary using the commutativity. Because deep learning is a non-linear transform, this model can include non-trivial interactions if more physical constraints are added to the proposed model.
4 Discussion
In the proposed model, it was essential to assume that a time evolution law exists for many-body system, as in cellular automata. Indeed, this rule is not as simple as block CA and can not be expressed in a mathematically closed form. Therefore, if this rule exists, how to find it is a difficult problem. But again, we can expect that this rule can also be found using deep learning algorithm. First of all, we have a way to test a system in the standard model. Therefore, if datasets are collected by conducting numerous experiments, the evolution law will be able to be learned. In fact, if we assume that random data really evolves according to the block CA rule as an example in the previous section, it has been confirmed that the block CA rule can be learned through an appropriate deep learning architecture and numerous random dataset. Moreover, it can be confirmed that learning is possible in reverse time as well.
Meanwhile, the proposed model does not represent the real world because it is made up of minimum simple assumptions. In order to propose a model that is close to reality, more physical axioms should be reflected in the model. However, the advantage of deep learning algorithms is the fact that they are flexible in constructing the architecture. For example, in order for the operator to satisfy covariance, the weight sharing condition should be enforced in the proposed deep learning architecture[10]. In addition, the activation function must be carefully selected to represent the CNN in covariant form. For the case of G-CNN, activation function satisfies covariance when a pointwise function is used for activation. In this way, it can be expected that the more physical requirements are mapped to the components that make up deep learning architecture, the closer the model will be to reality.